# Listen to Look into the Future: Audio-Visual Egocentric Gaze Anticipation

### 07 Dec 2023 v2

@@ 5 Conclusion
In this paper, we propose a novel contrastive spatial-temporal separable fusion approach (CSTS) for egocentric gaze anticipation. Our key contribution is to break down the fusion of the audio and visual modalities into a separate spatial fusion module for learning the spatial co-occurrence of salient visual features and audio signals, and a temporal fusion module for modeling the audio-visual correlations across different time steps. We further adopt a contrastive loss on the reweighted audio-visual representations from the fusion modules to facilitate multimodal representation learning. We demonstrate the benefits of our proposed model design on two egocentric video datasets: Ego4D and Aria. Our work is a key step for probing into human cognitive process with computational models, and provides important insights into multimodal representation learning, visual forecasting, and egocentric video understanding.
@@
This is the supplementary material for the paper titled ”Listen to Look into the Future: Audio-Visual Egocentric Gaze Anticipation”. We organize the content as follows:
  • A – Implementation Details
  • B – Details of Baseline Fusion Strategies
  • C – Experiments on Contrastive Learning
  • D – Results on Egocentric Gaze Estimation
  • E – Additional Visualization
  • F – Limitation and Future Work
  • G – Code and License
Appendix A Implementation Details
Architecture. Inspired by [14], we use a light-weight audio encoder composed of four self-attention blocks from MViT [12]. The model architecture is further detailed in Table 7. We initialize the video encoder with Kinetics-400 pretraining [30] and initialize the audio encoder using Xavier initialization [15]. The resulting video embeddings $\phi(x)$ have a dimension of $T=4,H=8,W=8,D=768$, and the resulting audio embeddings $\psi(a)$ have a dimension of $T=4,M=64,D=768$. We follow [36] to map audio-visual representation vectors to dimension $D^{\prime}=256$ for the contrastive loss. The output from the decoder is a downsampled heatmap which is upsampled to match the input size using trilinear interpolation. Following [33], we add intermediate features from each video encoder block to the corresponding decoder block output via skip connections to compensate for the loss of low-level textures.
Appendix B Details of Baseline Fusion Strategies
We compare with multiple different audio-visual fusion strategies in main paper Table 2. The details of each baseline are listed as follows:
Seq. ST. We implement a sequential variant of our proposed fusion strategy. The visual and audio embeddings are fed into the temporal fusion module first and then go through the spatial fusion module. Specifically, we input $\phi(x)$ and $\psi(a)$ into the temporal fusion module and obtain the output $u_{v,t}$ and $u_{a,t}$. We also use $u_{a,t}$ to reweight $\psi(a)$ resulting in $u_{a}\in\mathbb{R}^{T\times M\times D}$ as discussed in Section 3.2 and 3.3 from the main paper. For the visual embeddings, we use $u_{v,t}$ to reweight $\phi(x)$ and obtain a new reweighted visual embedding, i.e. $\hat{u}_{v}=\phi(x)\otimes u_{v,t}\in\mathbb{R}^{T\times H\times W\times D}$. $\hat{u}_{v}$ and $u_{a}$ serve as the input to the spatial fusion module. Then, we further generate a new visual representation incorporating audio information from the spatial fusion module and input the resulting embedding into the decoder.
Appendix C Experiments on Contrastive Learning
@@
F1 Rec. Prec. F1 Rec. Prec.
STS + Vanilla Contr 39.0 53.7 30.6 59.1 66.5 53.1
STS + S-Contr 38.5 53.5 30.0 59.0 66.3 53.1
STS + T-Contr 38.9 54.0 30.5 59.0 66.7 53.0
STS + Cross Contr 38.9 54.4 30.2 59.3 66.8 53.3
STS + Post Contr 39.7 53.3 31.6 59.9 66.8 54.3
Table 5: Study of different strategies for contrastive loss implementation. The best results are highlighted with boldface. The highlighted row refers to the complete CSTS model. See Section C for further discussion.
In our model, we propose to use the audio-visual representations obtained after fusion (i.e. $u_{v}$ and $u_{a}$) to calculate contrastive loss. As a comparison, we also implement a baseline by feeding the raw embeddings from the encoders (i.e. $\phi(x)$ and $\psi(a)$) to the contrastive loss which is denoted as Vanilla Contr. To further investigate the contribution of contrastive learning, we also conduct experiments with three additional strategies:
We implement all contrastive learning baselines above on our proposed model architecture and fusion strategy (i.e. STS). The results are summarized in Table 5. Both S-Contr and T-Contr lag behind or perform on par with Vanilla Contr. One possible reason is that conducting contrastive learning using features obtained from only one fusion branch may compromise the representation learning of the other branch. Additionally, Cross Contr works on-par with Vanilla Contr on Ego4D but performs better on Aria. It also consistently outperforms S-Contr and T-Contr. This result validates our claim that implementing contrastive loss with reweighted representations from both spatial and temporal fusion leads to more gains for egocentric gaze anticipation. Moreover, our proposed strategy (reweighting the raw audio embedding $\psi(a)$ rather than the fused embedding after spatial fusion) outperforms Cross Contr. This is because in Cross Contr $u_{a,s}$ is derived from spatial fusion, where each audio token is fused with 64 visual tokens in the spatial fusion branch resulting in the dilution of audio features. All results further demonstrate the benefits of our proposed contrastive learning strategy.
Appendix D Results on Egocentric Gaze Estimation
In addition to egocentric gaze anticipation, we also evaluate the generalization capability of our proposed method by conducting experiments on the separate egocentric gaze estimation problem using our model design. We use the same experiment setup from the recent state-of-the-art method [33] in egocentric gaze estimation.
Methods Ego4D Aria
F1 Rec. Prec. F1 Rec. Prec.
@@ Appendix D Results on Egocentric Gaze Estimation
Table 6: Comparison with previous state-of-the-art models on egocentric gaze estimation. The best results are highlighted with boldface. See Section D for further discussion.
As demonstrated in Table 6, previous work [33] has shown the superiority of using a transformer-based architecture for egocentric gaze estimation. By incorporating the audio modality, CSTS surpasses the backbone MViT [12] (vision-only counterpart) by +2.8% on both Ego4D on Aria in terms of F1. These results indicate the audio modality makes important contributions to the performance on egocentric gaze estimation. Furthermore, our model outperforms GLC [33] by +0.6% and +1.3% on Ego4D and Aria respectively, achieving a new state-of-the-art performance for this problem. However, our method has a smaller performance improvement on the gaze estimation task compared to gaze anticipation. We speculate that this is because a person typically orients their gaze in response to audio signals, so there is usually a delay between audio and correlated gaze movement. The audio thus has a stronger connection with future gaze targets.
Appendix E Additional Visualization
@@
We showcase further qualitative comparisons with all the baselines for egocentric gaze anticipation in Fig. 6. We observe CSTS makes the most accurate predictions. We also illustrate some typical failure cases in Fig. 7. In the first example, our model makes an accurate prediction in the first frame but fails at the following time steps due to the gaze movement. In the second example, the the camera view and gaze target move from the left to the right. This drastic change causes the mistake in our model’s predictions. Similar failures also happen in the predictions of all baselines. Notably, existing deep models tend to only successfully anticipate steady gaze fixations or small gaze movements in the near future, and can not effectively capture large gaze shifts. This is the a common limitation shared by many existing works of future anticipation [28] in egocentric videos.
Appendix F Limitation and Future Work
@@
Appendix G Code and License
The usage of the Aria Dataset is under the Apache 2.0 License [^2] https://github.com/facebookresearch/vrs/blob/main/LICENSE, and the usage of the Ego4D Dataset is under the license agreement [^3] https://ego4d-data.org/pdfs/Ego4D-Licenses-Draft.pdf. Our implementation is built on top of [11], which is under the Apache License [^4] https://github.com/facebookresearch/SlowFast/blob/main/LICENSE. Our code and the train/test split on Aria dataset will be available at: https://www.anonymous-for-review.com.

### 22 Mar 2024 v3

@@ 5 Conclusion
In this paper, we propose a novel contrastive spatial-temporal separable fusion approach (CSTS) for egocentric gaze anticipation. Our key contribution is breaking down the fusion of the audio and visual modalities into a separate spatial fusion module for learning the spatial co-occurrence of visual features and audio signals, and a temporal fusion module for modeling the changing viewpoint and scene driven by audio stimuli. We further adopt a contrastive loss on the reweighted audio-visual representations from the fusion modules to facilitate multimodal representation learning. We demonstrate the benefits of our proposed model design on two egocentric video datasets: Ego4D and Aria. Our work is a key step for probing into human cognitive process with computational models, and provides important insights into multimodal representation learning, visual forecasting and egocentric video understanding.
@@
This is the supplementary material for the paper titled "Listen to Look into the Future: Audio-Visual Egocentric Gaze Anticipation". We organize the content as follows:
  • $\bullet$
    A – Comparison with Prior Audio-Visual Learning Strategies
  • $\bullet$
    B – Additional Experiment Results
    • $\diamond$
      B.1 – Experiments about Model Generalization Capability
    • $\diamond$
      B.2 – Experiments on Egocentric Gaze Estimation
    • $\diamond$
      B.3 – Additional Experiments on Contrastive Learning
    • $\diamond$
      B.4 – Additional Visualization
  • $\bullet$
    C – More Implementation Details
    • $\diamond$
      C.1 – Implementation Details of Our Model
    • $\diamond$
      C.2 – Implementation Details of Baseline Fusion Strategies
  • $\bullet$
    D – Limitation and Future Work
  • $\bullet$
    E – Code and License
Appendix A Comparison with Prior Audio-Visual Learning Strategies
We have specified the key differences between the egocentric action anticipation task and saliency prediction task in the second paragraph of Sec. 2 in the main paper. The experiment results also validate that our proposed spatial-temporal separable fusion strategy performs better in our task than other fusion strategies designed for saliency prediction and action recognition (please refer to Tab. 2 in the main paper). In this section, we further compare our model with typical audio-visual learning methods for saliency prediction and recognition tasks in terms of model design.
In Tab. 5, all prior methods are designed for exocentric videos (i.e., third-person videos) that have a fixed camera viewpoint through all frames. Though various fusion approaches are used in these methods, they fuse audio-visual embeddings jointly in time and space. In contrast, the egocentric gaze anticipation task has the unique challenges of moving viewpoint together the latency between the audio stimuli and human reactions. To address these challenges, our model uses a novel spatial-temporal separable fusion strategy which has not been studied in prior work. The experiments in Tab. 2 of the main paper shows that our method achieves the best performance in egocentric gaze anticipation task compared with prior audio-visual learning strategies. In addition, using contrastive learning to boost audio-visual representations in a specific task is still an understudied area. Huang et al. [27] use inter- and intra- contrastive loss to learn aligned audio and visual embeddings. However, they straightforwardly apply contrastive loss on the raw embeddings right after the encoders. In our model, we innovatively propose to adopt contrastive loss on the embeddings after fusion layers (i.e., post-fusion contrastive learning). We also validate its advantage in Tab. 3 of the main paper. These key differences consolidate our contributions and clearly distinguish our model from other audio-visual learning methods.
Table 5: Comparison with typical audio-visual learning methods for audio-visual saliency prediction and recognition. If more than one fusion strategies have been tried in one method, we only show the strategy leading to the best performance.
Methods View Fusion Cntr Architecture Task
Tavakoli et al. [64] Exo Concatenation w/o CNN Saliency Prediction
Min et al. [47] Exo Correlation Analysis w/o CNN Saliency Prediction
Tsiami et al. [67] Exo Bilinear w/o CNN Saliency Prediction
Yao et al. [75] Exo Inner Product w/o CNN Saliency Prediction
Change et al. [6] Exo Bilinear w/o CNN Saliency Prediction
Jain et al. [31] Exo Bilinear w/o CNN Saliency Prediction
Wang et al. [68] Exo Concatenation w/o CNN Saliency Prediction
Xiong et al. [73] Exo Self-Attentation w/o CNN Saliency Prediction
Nagrani et al. [51] Exo Attention Bottleneck w/o Transformer Video Classification
Huang et al. [27] Exo Self-Attention w/ Transformer Video Classification
Gao et al. [16] Exo Linear w/o LSTM Action Recognition
Kazakos et al. [34] Exo Linear w/o CNN Action Recognition
Wang et al. [70] Exo Weigted Sum w/o CNN Action Recognition
Xiao et al. [72] Exo Self-Attentation w/o CNN Action Recognition
Liu et al. [41] Exo Linear w/o CNN Action Recognition
Senocak et al. [59] Exo Linear w/o CNN Action Recognition
Praveen et al. [53] Exo Self-Attention w/o CNN Emotion Recognition
Chudasama et al. [9] Exo Self-Attention w/o Transformer Emotion Recognition
CSTS (Ours) Ego Spatial-Temporal w/ Transformer Gaze Anticipation
Separable
Appendix B Additional Experiment Results
Table 6: Zero-shot experiments on Aria dataset. All baselines and our model are trained only on Ego4D training set. We consider F1 score as the primary metric in our experiments. The green row refers to our model, and the best results are highlighted with boldface. See Sec. B.1 for further discussion.
Methods F1 Score Recall Precision
GazeMLE [38] 44.0 59.0 35.0
AttnTransit [29] 43.1 57.5 34.5
I3D-R50 [15] 41.5 77.2 28.4
MViT [14] 44.1 59.7 35.0
GLC [36] 46.9 72.8 34.6
[…]
DFG+ [76] 43.1 76.4 30.0
CSTS 50.8 62.2 42.9
B.1 Experiments about Model Generalization Capability
To validate the generalization capability of our model, we compare our model with prior state-of-the-art models in a zero-shot setting. Specifically, We train our model and all baselines with Ego4D training set and test them with Aria test set. Note that the Aria data is invisible to all models during training. The results are presented in Tab. 6. Our model outperforms the best egocentric gaze anticipation model (DFG+) by +7.7% and also exceeds the strongest baseline (GLC) by +3.9% in F1 score (primary metric). The remarkable improvement suggests that, with our novel fusion and contrastive learning approaches, our model is able to generalize better to other unseen data, which is critical for applying it to real-world problems.
Table 7: Comparison with prior state-of-the-art models on egocentric gaze estimation. The green row refers to our model. The best results are highlighted with boldface. See Sec. B.2 for further discussion.
@@ B.1 Experiments about Model Generalization Capability
F1 Score Recall Precision F1 Score Recall Precision
@@
B.2 Experiments on Egocentric Gaze Estimation
In addition to egocentric gaze anticipation, we also evaluate the advantage of our model in another gaze modeling problem – egocentric gaze estimation. Instead of forecasting future gaze, egocentric gaze estimation requires gaze target prediction in the current video frames. We use the same experiment setup from the recent state-of-the-art method [36].
As demonstrated in Tab. 7, the prior work [36] has shown the superiority of using a transformer-based architecture for egocentric gaze estimation. By incorporating the audio modality, CSTS surpasses the backbone MViT [14] (vision-only counterpart) by +2.8% on both Ego4D on Aria in terms of F1 score. These results indicate the audio modality also makes important contributions to the performance on egocentric gaze estimation. Furthermore, our model outperforms GLC [36] by +0.6% and +1.3% on Ego4D and Aria respectively, achieving a new state-of-the-art performance for this problem. However, our method has a smaller performance improvement on the gaze estimation task compared to gaze anticipation. The possible reason is that the audio stream has a stronger connection with future gaze targets than current gaze behaviors because of the natural latency between the audio stimuli and human reactions.
B.3 Additional Experiments on Contrastive Learning
Table 8: Study of different strategies for contrastive loss implementation. Post Cntr refers to our proposed post-fusion contrastive learning strategy, and the green row refers to the complete CSTS model. The best results are highlighted with boldface. See Sec. B.3 for further discussion.
Methods Ego4D Aria
F1 Score Recall Precision F1 Score Recall Precision
STS + Vanilla Contr 39.0 53.7 30.6 59.1 66.5 53.1
STS + S-Contr 38.5 53.5 30.0 59.0 66.3 53.1
STS + T-Contr 38.9 54.0 30.5 59.0 66.7 53.0
STS + Cross Contr 38.9 54.4 30.2 59.3 66.8 53.3
STS + Post Contr 39.7 53.3 31.6 59.9 66.8 54.3
In our model, we propose to use the audio-visual representations obtained after fusion (i.e. $u_{v}$ and $u_{a}$) to calculate contrastive loss (i.e., post-fusion contrastive learning). As a comparison, we also implement a baseline by feeding the raw embeddings from the encoders (i.e. $\phi(x)$ and $\psi(a)$) to the contrastive loss which is denoted as Vanilla Contr. To further investigate the contribution of contrastive learning, we also conduct experiments with three additional strategies:
We implement all contrastive learning baselines above on our proposed model architecture and fusion strategy (i.e., STS). The results are summarized in Tab. 8. Both S-Contr and T-Contr lag behind or perform on par with Vanilla Contr. One possible reason is that conducting contrastive learning using features obtained from only one fusion branch may compromise the representation learning of the other branch. Additionally, Cross Contr works on-par with Vanilla Contr on Ego4D but performs better on Aria. It also consistently outperforms S-Contr and T-Contr. This result validates our claim that implementing contrastive loss with reweighted representations from both spatial and temporal fusion leads to more gains for egocentric gaze anticipation. Moreover, our proposed strategy (reweighting the raw audio embedding $\psi(a)$ rather than the fused embedding after spatial fusion) outperforms Cross Contr. This is because in Cross Contr $u_{a,s}$ is derived from spatial fusion, where each audio token is fused with 64 visual tokens in the spatial fusion branch resulting in the dilution of audio features. All results further demonstrate the benefits of our proposed contrastive learning strategy.
@@ B.3 Additional Experiments on Contrastive Learning
B.4 Additional Visualization
We showcase more qualitative comparisons with all the baselines for egocentric gaze anticipation in Fig. 6. We observe CSTS makes the most accurate predictions. We also illustrate some typical failure cases in Fig. 7. In the first example, our model makes an accurate prediction in the first frame but fails at the following time steps due to the gaze movement. In the second example, the the camera view and gaze target move from the left to the right. This drastic change causes the mistake in our model’s predictions. Similar failures also happen in the predictions of all baselines. Notably, existing deep models tend to only successfully anticipate steady gaze fixations or small gaze movements in the near future, and can not effectively capture large gaze shifts. This is the a common limitation shared by many existing works of future anticipation [32] in egocentric videos.
Appendix C More Implementation Details
C.1 Implementation Details of Our Model
Architecture. Inspired by [17], we use a light-weight audio encoder composed of four self-attention blocks from MViT [14]. The model architecture is further detailed in Tab. 9. We initialize the video encoder with Kinetics-400 pretraining [33] and initialize the audio encoder using Xavier initialization [18]. The resulting video embeddings $\phi(x)$ have a dimension of $T=4,H=8,W=8,D=768$, and the resulting audio embeddings $\psi(a)$ have a dimension of $T=4,M=64,D=768$. We follow [39] to map audio-visual representation vectors to dimension $D^{\prime}=256$ for the contrastive loss. The output from the decoder is a downsampled heatmap which is upsampled to match the input size using trilinear interpolation. Following [36], we add intermediate features from each video encoder block to the corresponding decoder block output via skip connections to compensate for the loss of low-level textures.
C.2 Implementation Details of Baseline Fusion Strategies
We compare with multiple different audio-visual fusion strategies in main paper Tab. 2. The details of each baseline are listed as follows:
Linear. We reshape the video embedding and audio embedding to the shape $\hat{N}\times D$. We concatenate the two reshaped embeddings (resulting in the dimension of $\hat{N}\times 2D$) and input it to two linear layers. The dimension of the output is $\hat{N}\times D$ and we reshape it back to $T\times H\times W\times D$ which is fed into the decoder.
Bilinear. We reduce the length of video tokens and audio tokens to 256 using a linear layer respectively. Then we input the resulting video and audio tokens into a bilinear layer. The output is fed into the decoder for gaze forecasting.
Appendix D Limitation and Future Work
@@
Appendix E Code and License
[…]